OpenStudy (anonymous):

Please help! Factor out the GCF from the polynomial: 7x+14=?

jimthompson5910 (jim_thompson5910):

7x+14 is the same as 7*x+7*2

jimthompson5910 (jim_thompson5910):

now use the distributive property

OpenStudy (anonymous):

Would it be 7(x+7)

jimthompson5910 (jim_thompson5910):

close, but no

OpenStudy (anonymous):

I'm so lost!

jimthompson5910 (jim_thompson5910):

hint: a*(b+c) = a*b + a*c that is the distributive property

OpenStudy (anonymous):

7(x+7)= 7x+14????

jimthompson5910 (jim_thompson5910):

7(x+7) = 7*x + 7*7 7(x+7) = 7x + 49 so that's not 7x + 14 like we want

OpenStudy (anonymous):

So is it 7(x+7) = 7x + 49????

jimthompson5910 (jim_thompson5910):

the original problem is 7x+14

jimthompson5910 (jim_thompson5910):

not 7x+49

jimthompson5910 (jim_thompson5910):

you need to factor 7x + 14 NOT 7x + 49